# Group-robust Machine Unlearning

### v1

@@ 5 Conclusion
| L1-sparse [23] | 99.00.1 | 60.016.3 | 85.30.8 | 46.724.9 | 29.21.7 | 57.92.6 | 92.33.0 | |
| SalUn [13] | 100.00.0 | 53.39.4 | 78.33.3 | 66.79.4 | 42.75.1 | 33.37.3 | 81.66.7 | |
| SCRUB [28] | 98.90.1 | 53.39.4 | 86.90.4 | 53.318.9 | 31.11.1 | 44.33.5 | 91.32.0 | |
@@
| L1-sparse [23] | 99.00.1 | 73.318.9 | 85.31.2 | 53.324.9 | 29.12.8 | 58.23.8 | 92.23.0 | |
| SalUn [13] | 99.90.1 | 53.39.4 | 76.72.9 | 86.79.4 | 45.94.7 | 29.67.4 | 81.32.9 | |
| SCRUB [28] | 98.80.2 | 53.39.4 | 86.80.7 | 60.016.3 | 31.50.8 | 42.83.5 | 89.80.4 | |
@@
| L1-sparse [23] | 99.00.1 | 59.58.9 | 85.60.4 | 44.011.8 | 32.21.8 | 48.87.4 | 94.40.3 | |
| SalUn [13] | 100.00.0 | 50.05.1 | 81.80.4 | 90.53.4 | 38.71.2 | 39.33.3 | 87.43.5 | |
| SCRUB [28] | 98.80.2 | 60.77.7 | 86.90.6 | 45.28.9 | 31.91.8 | 41.71.7 | 93.91.4 | |
@@
| L1-sparse [23] | 98.70.1 | 64.35.8 | 85.01.2 | 46.412.7 | 30.61.4 | 53.74.3 | 94.71.1 | |
| SalUn [13] | 100.00.0 | 47.64.5 | 81.11.9 | 91.77.3 | 39.02.2 | 39.01.6 | 85.84.2 | |
| SCRUB [28] | 98.90.2 | 66.71.7 | 87.00.5 | 44.08.9 | 30.91.3 | 44.33.1 | 94.51.3 | |
@@
| L1-sparse [23] | 98.90.2 | 60.03.3 | 82.91.2 | 50.75.0 | 35.30.4 | 49.94.6 | 92.91.0 | |
| SalUn [13] | 100.00.0 | 40.04.3 | 81.30.9 | 92.73.4 | 41.41.3 | 30.82.2 | 89.62.0 | |
| SCRUB [28] | 97.80.1 | 30.71.9 | 86.10.5 | 52.73.4 | 36.61.0 | 25.11.5 | 92.70.7 | |
@@
| L1-sparse [23] | 99.00.2 | 59.310.6 | 84.60.6 | 45.36.8 | 31.50.7 | 55.04.0 | 91.54.1 | |
| SalUn [13] | 100.00.0 | 45.32.5 | 80.30.7 | 87.31.9 | 41.80.5 | 31.94.8 | 90.91.0 | |
| SCRUB [28] | 98.00.1 | 33.33.4 | 86.20.7 | 54.75.2 | 35.91.5 | 28.03.4 | 93.61.2 | |
@@
| L1-sparse [23] | 63.70.3 | 78.93.5 | 56.10.8 | 0.00.0 | 5.52.6 | 69.72.2 | 97.30.7 | |
| SalUn [13] | 65.90.8 | 73.93.9 | 55.11.1 | 0.50.0 | 2.91.1 | 69.87.0 | 97.80.9 | |
| SCRUB [28] | 68.40.5 | 78.70.8 | 57.50.3 | 0.20.2 | 5.70.2 | 70.41.5 | 97.90.5 | |
@@
| L1-sparse [23] | 64.00.3 | 72.70.7 | 56.40.6 | 0.20.2 | 5.31.2 | 69.10.7 | 98.60.3 | |
| SalUn [13] | 66.20.4 | 80.11.5 | 55.30.4 | 0.20.2 | 4.70.9 | 73.34.6 | 97.10.3 | |
| SCRUB [28] | 68.40.5 | 79.21.0 | 57.50.4 | 0.20.2 | 5.61.1 | 70.91.7 | 97.80.6 | |
@@
| L1-sparse [23] | 64.00.3 | 74.11.2 | 56.90.5 | 0.20.1 | 6.10.7 | 69.40.7 | 98.30.2 | |
| SalUn [13] | 66.30.4 | 66.63.4 | 55.90.6 | 0.30.1 | 9.00.5 | 60.32.4 | 97.11.1 | |
| SCRUB [28] | 66.90.1 | 65.41.6 | 56.70.7 | 1.00.0 | 9.91.3 | 61.32.5 | 97.00.5 | |
@@
| L1-sparse [23] | 64.40.1 | 72.92.1 | 56.00.9 | 0.30.1 | 6.12.1 | 67.06.8 | 97.30.3 | |
| SalUn [13] | 65.10.4 | 69.86.3 | 54.80.6 | 0.30.2 | 6.62.1 | 63.73.4 | 97.20.4 | |
| SCRUB [28] | 66.70.1 | 73.42.2 | 57.20.5 | 0.70.3 | 6.21.1 | 70.22.7 | 98.70.7 | |
@@
| L1-sparse [23] | 64.50.2 | 57.12.1 | 55.20.6 | 0.40.1 | 13.00.7 | 51.01.9 | 97.70.3 | |
| SalUn [13] | 65.70.5 | 46.56.2 | 53.90.1 | 0.50.1 | 15.31.0 | 42.84.9 | 95.21.8 | |
| SCRUB [28] | 60.21.0 | 52.74.4 | 53.30.5 | 2.40.7 | 15.61.4 | 48.74.9 | 95.70.6 | |
@@
| L1-sparse [23] | 64.00.4 | 74.53.0 | 55.90.4 | 0.30.2 | 5.60.6 | 69.84.6 | 91.91.5 | |
| SalUn [13] | 65.50.5 | 66.13.7 | 55.30.2 | 0.50.3 | 7.21.2 | 62.84.9 | 94.91.7 | |
| SCRUB [28] | 61.21.1 | 65.53.5 | 54.50.3 | 1.30.1 | 9.50.9 | 64.42.7 | 94.51.6 | |
@@
| Method | reweight | RA | UA | TA | MIA | EO | GA | Avg. Gap |
@@
| L1-sparse [23] | 63.50.6 | 69.93.9 | 55.30.1 | 0.50.0 | 2.61.0 | 64.44.4 | 97.71.1 | |
| SalUn [13] | 64.30.5 | 64.61.2 | 54.00.1 | 0.20.1 | 3.60.3 | 59.71.5 | 96.00.6 | |
| SCRUB [28] | 67.20.4 | 74.30.7 | 56.90.2 | 0.30.1 | 1.80.7 | 65.60.6 | 97.70.3 | |
@@
| L1-sparse [23] | 63.70.1 | 75.20.6 | 56.30.1 | 0.20.1 | 1.30.3 | 69.61.2 | 98.60.2 | |
| SalUn [13] | 63.70.8 | 74.41.5 | 55.50.4 | 0.40.0 | 2.30.3 | 69.01.6 | 98.40.2 | |
| SCRUB [28] | 66.70.4 | 80.50.1 | 57.40.5 | 0.40.1 | 1.70.3 | 71.30.4 | 97.40.3 | |
@@
| L1-sparse [23] | 65.50.8 | 61.80.3 | 55.50.5 | 0.50.1 | 1.60.8 | 57.00.9 | 98.70.2 | |
| SalUn [13] | 64.50.5 | 60.03.8 | 55.21.0 | 0.50.1 | 1.10.4 | 57.03.8 | 98.00.7 | |
| SCRUB [28] | 66.70.6 | 62.41.0 | 55.40.3 | 0.30.1 | 1.70.9 | 55.20.6 | 98.40.2 | |
@@
| L1-sparse [23] | 64.30.7 | 66.70.4 | 55.80.4 | 0.40.1 | 2.10.8 | 60.81.0 | 98.30.3 | |
| SalUn [13] | 62.51.6 | 64.62.0 | 54.60.8 | 0.30.1 | 2.20.4 | 60.12.0 | 98.00.3 | |
| SCRUB [28] | 65.30.3 | 71.40.6 | 57.00.3 | 0.20.0 | 1.90.9 | 63.70.7 | 97.10.1 | |
@@
This supplementary material details the main paper’s experimental protocol and additional results. Appendix A discusses the implementation details, reporting MIU pseudocode, the hardware used, training hyperparameters, and the implementation of chosen baselines. Then, Appendix B formalizes the metrics used, Appendix C reports results that could not be included in the main paper, and Appendix D describes our work’s limitations and potential negative societal impacts.
Appendix A Implementation details
This section presents a general overview of the implementation details of this paper, by highlighting MIU pseudocode, the hardware used to perform the experiments, the software, and the algorithms of choice.
MIU pseudocode. Algorithm 1 presents MIU pseudocode in a PyTorch-like [39] style. We follow previous works [28, 13] and compute unlearning and retaining steps separately. Thus, we alternate between computing an unlearning epoch using Eq. 3 and a retaining one using Eqs. 5 and 4. Like SCRUB [28], we find it beneficial to stop performing unlearning steps after a predefined number of epochs (i.e., 5 out of 10). Contrary to Ragonesi et al. [40], we do not update the mutual information representation at every step. Instead, we empirically observed that updating it via 100 gradient updates for the first epoch and 10 updates for the remaining iterations is sufficient to achieve satisfactory results and limit the required resources. As we keep the optimization steps number fixed, the overall mutual information estimation overhead depends on the dataset size. For a small dataset like Waterbirds [42], we estimate an overhead of about . Instead, for CelebA [31] and FairFace [24], we estimate an increase of approximately in unlearning time, which is negligible.
Hardware and training details. As Sec. 4.1 mentions, we compute Pretrain and Retrain by fine-tuning a ResNet-18 [17] pre-trained on ImageNet [41] for 30 epochs, using SGD with 0.9 momentum and weight decay. The learning rate is decayed with a cosine annealing scheduler for the entire training. We additionally warm-up the learning rate for the first two epochs using a linear scheduler. We apply standard data augmentation techniques such as random resized crop, random horizontal flip, and input normalization [17]. We limited the fine-tuning to 10 epochs for unlearning methods, searching for the optimal configuration for the other hyperparameters. The parameter of MIU is set between 1 and 10 (see Sec. C.4). We ran all experiments on a single A100 Nvidia GPU, using PyTorch [39].
Baselines. As Sec. 4.1 mentions, we reimplemented all three investigated baselines to experiment easily in group-robust unlearning following the existing codebases. For L1-sparse [23], we perform 10 fine-tuning epochs on the retain set with a linearly decaying L1 regularization that follows this rule: , where is the epoch, is the total number of iterations, and the initial penalty. For SalUn [13] we followed the “small-scale” implementation of the original codebase (i.e., the implementation for CIFAR-10 [27] and SVHN [34]) as it is meant for datasets with few classes. We first compute the saliency mask using the gradient information from the forget set, pruning 50% of the network weights. We tune the remaining weights for 10 epochs by alternating a full pass on the forget set and a full pass on the retain set. Also, SCRUB [28] is implemented by separating forget and retain steps, following the original code. The forget step maximizes the KL divergence between the original and the unlearned model in the forget data set. Following the original paper, we stop computing it after 3, 5, or 7 epochs. Instead, the retain step minimizes the linear combination between the cross-entropy loss and the KL divergence between the original and unlearned model, respectively scaled by 0.99 and 0.001, as reported in [28]. We note that all methods use the same dataset splits; therefore, they must unlearn the same forget set.
Appendix B Metrics
This section provides further details on the adopted metrics, i.e., the RA, UA, TA, the membership-inference attack (MIA) [47], the dominant forget group accuracy (GA), the equalized odds (EO), and the average gap (Avg. Gap) [13].
RA, UA, and TA. We evaluate the model on the retain, forget, and test sets to compute the retain, forget, and test accuracy. Therefore, we report the ratio of correctly classified samples for each of these subsets of the dataset.
MIA. To compute the membership inference attack, we follow previous works [23, 13] and train a model on retain and validation losses to predict membership. Therefore, we assign a different binary label to retain and validation losses and train the model to discriminate them. After model training, we use such a model to infer the membership of forget set data. In all tables, we report the MIA-Efficacy [23] that is computed as follows:
where are the true negatives, i.e., the number of samples the MIA predicted as non-members. Instead of training a support vector machine as in [23, 13], we used a random forest as accuracies are comparable with an SVM. Moreover, training is faster since it can be easily parallelized. The higher the MIA-Efficacy, the better the model privacy protection.
EO. Equalized Odds [16] measures model fairness or prediction dependencies on protected attributes. To compute the EO we measure model performance discrepancy by varying the sensitive attribute value and averaging over different target labels. Formally EO for a binary classification model is computed as follows [16]:
where are random variables describing model predictions, target attributes, and sensitive attributes. EO measures the absolute difference in outputting a positive prediction when the protected attribute equals 1 and 0, averaging over the two target attributes. In the FairFace [24] case, we set all classes and protected attributes that do not match those of the dominant group of the forget set as and , as FairFace counts more than two classes and more than two attributes. The lower the EO the better the model fairness.
GA. Short for Group Accuracy, GA measures the accuracy of the dominant group of the forget set. In other words, it measures the ratio of correctly classified test samples that belong to the same dominant group(s) of the forget set. Therefore, let be the subset of the test set composed of all samples of group , then GA is computed as follows:
Avg. Gap. Following previous works [23, 13], we compute the Avg. Gap to simplify the comparison among different methodologies. Avg. Gap is computed as the average metric discrepancy between the unlearned model and the retrained gold standard. Formally, let be the set of all metrics used in this paper, then Avg. Gap is computed as follows:
Appendix C Additional results
This section contains the additional results that could not be included in the main paper. Section C.1 intuitively shows the accuracy degradation caused by non-uniformly sampled forget sets in all three investigated datasets. Sections C.2 and C.3 report tables associated with experiments on different unlearning ratios and sampling the forget set from multiple groups. Section C.4 further expands MIU ablations. Finally, Section C.5 further evaluates MIU’s and reweight’s fairness and robustness preservation.
C.1 Unlearning non-uniformly sampled data
Figure 2 of the main paper shows how the forget set dominant group accuracy drops when increasing the unlearning ratio, limiting the analysis to CelebA [31] due to space constraints. This section reports the same experiment on all three investigated datasets for completeness.
Figure 7 shows accuracy variations of attractive males, waterbirds on land, and 20-29 years old afro-americans, when unlearned from respectively CelebA [31], Waterbirds [42], and FairFace [24] with different unlearning ratios: , , . While the accuracy drop is more evident on CelebA [31], it is also visible in the other two datasets. Interestingly, some unlearning methods are more robust to this performance drop than others (e.g., L1-sparse [23] scores a 49.9% in Waterbirds [42] with unlearning ratio 0.9 vs. 30.8% of SalUn [13], which is the second best).
Although some methods are more robust, we argue that the quality of the unlearning process influences the accuracy of the dominant group of the forget set. The less the forget set was unlearned, the more the performance retention. As an example, Sec. 4.2 highlights that L1-sparse [23] fails to effectively unlearn the forget set in the FairFace [24] experiment, showing higher UA and GA (74.1% and 69.4%) than the Retrain (57.8% and 58.7%). Thus, to better investigate the relationship between unlearning effectiveness and performance retention, Sec. C.2 reports tables associated with Fig. 7 using all investigated metrics.
C.2 Impact of different unlearning ratios
For completeness purposes, this section reports all tables associated with Figs. 2, 7, 5 and 4. For these experiments, the main paper summarizes the results to reduce the occupied space and simplify the interpretation (i.e., limiting the reported metrics to one). Therefore, Tabs. 5, 7, 8, 10, 11 and 13 present experiments in CelebA [31], Waterbirds [42], and FairFace [24] datasets, with unlearning ratios of , and . Additionally, Tabs. 6, 9 and 12 report the standard deviations of Tabs. 1, 2 and 3.
Overall, we notice that in CelebA [31], the higher the unlearning ratio, the lower the forget accuracy (from 63.1% to 31.6% using MIU), with the gap being reduced when reweight is included in the retaining step (around 64-68% with MIU). RA, TA, and MIA remain stable across different unlearning ratios. Instead, EO behaves similarly to GA and UA, with high values at high unlearning ratios (e.g., MIU scores 22.3% vs. 38.3% with 0.1 and 0.9 unlearning ratios).
In Waterbirds [42], metrics do not show global trends, except for EO and GA, which worsen as the unlearning ratio grows. Therefore, methods lose accuracy on the dominant group of the forget set as the ratio of unlearned samples grows, with SCRUB [28] showing the highest drop (from 44.3% to 25.1%). Similarly, GA and EO worsen even when adding reweight. As the unlearning ratio grows, the number of forget set dominant group samples left in the retain set lowers. Given the small size of Waterbirds [42] (4795 samples, of which 56 are in the smallest group), reweight strongly increases the sampling likelihood of a restricted number of samples to preserve original accuracy, causing overfitting. Thus, the overall benefit of reweight is reduced (GA increases by only 1.3 with Retrain).
Also FairFace [24] shows a general drop in UA, which grows when reweight is applied. EO and GA also behave like in CelebA [31], with an enhanced degradation at higher unlearning ratios. However, MIU shows good robustness even without reweight, scoring EO and GA that are close to Retrain+reweight, e.g., 10.4% vs. 11.9% in EO and 56.8% vs. 53.9 in GA with a 0.9 unlearning ratio (Tab. 13).
C.3 Multi-group unlearning

### v2

@@ B.1 Unlearning non-uniformly sampled data
| L1-sparse (Jia et al., 2023) | 99.00.1 | 60.016.3 | 85.30.8 | 46.724.9 | 29.21.7 | 57.92.6 | 92.33.0 | |
| SalUn (Fan et al., 2024b) | 100.00.0 | 53.39.4 | 78.33.3 | 66.79.4 | 42.75.1 | 33.37.3 | 81.66.7 | |
| SCRUB (Kurmanji et al., 2024) | 98.90.1 | 53.39.4 | 86.90.4 | 53.318.9 | 31.11.1 | 44.33.5 | 91.32.0 | |
@@
| L1-sparse (Jia et al., 2023) | 99.00.1 | 73.318.9 | 85.31.2 | 53.324.9 | 29.12.8 | 58.23.8 | 92.23.0 | |
| SalUn (Fan et al., 2024b) | 99.90.1 | 53.39.4 | 76.72.9 | 86.79.4 | 45.94.7 | 29.67.4 | 81.32.9 | |
| SCRUB (Kurmanji et al., 2024) | 98.80.2 | 53.39.4 | 86.80.7 | 60.016.3 | 31.50.8 | 42.83.5 | 89.80.4 | |
@@
| L1-sparse (Jia et al., 2023) | 99.00.1 | 59.58.9 | 85.60.4 | 44.011.8 | 32.21.8 | 48.87.4 | 94.40.3 | |
| SalUn (Fan et al., 2024b) | 100.00.0 | 50.05.1 | 81.80.4 | 90.53.4 | 38.71.2 | 39.33.3 | 87.43.5 | |
| SCRUB (Kurmanji et al., 2024) | 98.80.2 | 60.77.7 | 86.90.6 | 45.28.9 | 31.91.8 | 41.71.7 | 93.91.4 | |
@@
| L1-sparse (Jia et al., 2023) | 98.70.1 | 64.35.8 | 85.01.2 | 46.412.7 | 30.61.4 | 53.74.3 | 94.71.1 | |
| SalUn (Fan et al., 2024b) | 100.00.0 | 47.64.5 | 81.11.9 | 91.77.3 | 39.02.2 | 39.01.6 | 85.84.2 | |
| SCRUB (Kurmanji et al., 2024) | 98.90.2 | 66.71.7 | 87.00.5 | 44.08.9 | 30.91.3 | 44.33.1 | 94.51.3 | |
@@
| L1-sparse (Jia et al., 2023) | 98.90.2 | 60.03.3 | 82.91.2 | 50.75.0 | 35.30.4 | 49.94.6 | 92.91.0 | |
| SalUn (Fan et al., 2024b) | 100.00.0 | 40.04.3 | 81.30.9 | 92.73.4 | 41.41.3 | 30.82.2 | 89.62.0 | |
| SCRUB (Kurmanji et al., 2024) | 97.80.1 | 30.71.9 | 86.10.5 | 52.73.4 | 36.61.0 | 25.11.5 | 92.70.7 | |
@@
| L1-sparse (Jia et al., 2023) | 99.00.2 | 59.310.6 | 84.60.6 | 45.36.8 | 31.50.7 | 55.04.0 | 91.54.1 | |
| SalUn (Fan et al., 2024b) | 100.00.0 | 45.32.5 | 80.30.7 | 87.31.9 | 41.80.5 | 31.94.8 | 90.91.0 | |
| SCRUB (Kurmanji et al., 2024) | 98.00.1 | 33.33.4 | 86.20.7 | 54.75.2 | 35.91.5 | 28.03.4 | 93.61.2 | |
@@
| L1-sparse (Jia et al., 2023) | 63.70.3 | 78.93.5 | 56.10.8 | 0.00.0 | 5.52.6 | 69.72.2 | 97.30.7 | |
| SalUn (Fan et al., 2024b) | 65.90.8 | 73.93.9 | 55.11.1 | 0.50.0 | 2.91.1 | 69.87.0 | 97.80.9 | |
| SCRUB (Kurmanji et al., 2024) | 68.40.5 | 78.70.8 | 57.50.3 | 0.20.2 | 5.70.2 | 70.41.5 | 97.90.5 | |
@@
| L1-sparse (Jia et al., 2023) | 64.00.3 | 72.70.7 | 56.40.6 | 0.20.2 | 5.31.2 | 69.10.7 | 98.60.3 | |
| SalUn (Fan et al., 2024b) | 66.20.4 | 80.11.5 | 55.30.4 | 0.20.2 | 4.70.9 | 73.34.6 | 97.10.3 | |
| SCRUB (Kurmanji et al., 2024) | 68.40.5 | 79.21.0 | 57.50.4 | 0.20.2 | 5.61.1 | 70.91.7 | 97.80.6 | |
@@
| L1-sparse (Jia et al., 2023) | 64.00.3 | 74.11.2 | 56.90.5 | 0.20.1 | 6.10.7 | 69.40.7 | 98.30.2 | |
| SalUn (Fan et al., 2024b) | 66.30.4 | 66.63.4 | 55.90.6 | 0.30.1 | 9.00.5 | 60.32.4 | 97.11.1 | |
| SCRUB (Kurmanji et al., 2024) | 66.90.1 | 65.41.6 | 56.70.7 | 1.00.0 | 9.91.3 | 61.32.5 | 97.00.5 | |
@@
| L1-sparse (Jia et al., 2023) | 64.40.1 | 72.92.1 | 56.00.9 | 0.30.1 | 6.12.1 | 67.06.8 | 97.30.3 | |
| SalUn (Fan et al., 2024b) | 65.10.4 | 69.86.3 | 54.80.6 | 0.30.2 | 6.62.1 | 63.73.4 | 97.20.4 | |
| SCRUB (Kurmanji et al., 2024) | 66.70.1 | 73.42.2 | 57.20.5 | 0.70.3 | 6.21.1 | 70.22.7 | 98.70.7 | |
@@
| L1-sparse (Jia et al., 2023) | 64.50.2 | 57.12.1 | 55.20.6 | 0.40.1 | 13.00.7 | 51.01.9 | 97.70.3 | |
| SalUn (Fan et al., 2024b) | 65.70.5 | 46.56.2 | 53.90.1 | 0.50.1 | 15.31.0 | 42.84.9 | 95.21.8 | |
| SCRUB (Kurmanji et al., 2024) | 60.21.0 | 52.74.4 | 53.30.5 | 2.40.7 | 15.61.4 | 48.74.9 | 95.70.6 | |
@@
| L1-sparse (Jia et al., 2023) | 64.00.4 | 74.53.0 | 55.90.4 | 0.30.2 | 5.60.6 | 69.84.6 | 91.91.5 | |
| SalUn (Fan et al., 2024b) | 65.50.5 | 66.13.7 | 55.30.2 | 0.50.3 | 7.21.2 | 62.84.9 | 94.91.7 | |
| SCRUB (Kurmanji et al., 2024) | 61.21.1 | 65.53.5 | 54.50.3 | 1.30.1 | 9.50.9 | 64.42.7 | 94.51.6 | |
@@
B.2 Impact of different unlearning ratios
For completeness purposes, this section reports all tables associated with Figs.˜2, 7, 6 and 4. For these experiments, the main paper summarizes the results to reduce the occupied space and simplify the interpretation (i.e., limiting the reported metrics to one). Therefore, Tabs.˜5, 7, 8, 10, 11 and 13 present experiments in CelebA (Liu et al., 2015), Waterbirds (Sagawa et al., 2020), and FairFace (Karkkainen & Joo, 2021) datasets, with unlearning ratios of , and . Additionally, Tabs.˜6, 9 and 12 report the standard deviations of Tabs.˜1, 2 and 3.
Overall, we notice that in CelebA (Liu et al., 2015), the higher the unlearning ratio, the lower the forget accuracy (from 43.0% to 32.6% using MIU), with the gap being reduced when reweight is included in the retaining step (from 44.9% to 35.2% with MIU). RA, TA, and MIA remain stable across different unlearning ratios. Instead, EO behaves similarly to GA and UA, with high values at high unlearning ratios (e.g., MIU scores 22.9% vs. 26.9% with 0.1 and 0.9 unlearning ratios).
In Waterbirds (Sagawa et al., 2020), metrics do not show global trends, except for EO and GA, which worsen as the unlearning ratio grows. Therefore, methods lose accuracy on the dominant group of the forget set as the ratio of unlearned samples grows, with SCRUB (Kurmanji et al., 2024) showing the highest drop (from 44.3% to 25.1%). Similarly, GA and EO worsen even when adding reweight. As the unlearning ratio grows, the number of forget set dominant group samples left in the remaining set lowers. Given the small size of Waterbirds (Sagawa et al., 2020) (4795 samples, of which 56 are in the smallest group), reweight strongly increases the sampling likelihood of a restricted number of samples to preserve original accuracy, causing overfitting. Thus, the overall benefit of reweight is reduced (GA increases by only 1.3 with Retrain).
Also FairFace (Karkkainen & Joo, 2021) shows a general drop in UA, which grows when reweight is applied. EO and GA also behave like in CelebA (Liu et al., 2015), with an enhanced degradation at higher unlearning ratios. However, MIU shows good robustness even without reweight, scoring EO and GA that are close to Retrain+reweight, e.g., 10.4% vs. 11.9% in EO and 56.8% vs. 53.9 in GA with a 0.9 unlearning ratio (Tab.˜13).
| method | RW | RA | UA | TA | MIA | EO | GA | avg. gap |
@@ B.2 Impact of different unlearning ratios
| L1-sparse (Jia et al., 2023) | 63.50.6 | 69.93.9 | 55.30.1 | 0.50.0 | 2.61.0 | 64.44.4 | 97.71.1 | |
| SalUn (Fan et al., 2024b) | 64.30.5 | 64.61.2 | 54.00.1 | 0.20.1 | 3.60.3 | 59.71.5 | 96.00.6 | |
| SCRUB (Kurmanji et al., 2024) | 67.20.4 | 74.30.7 | 56.90.2 | 0.30.1 | 1.80.7 | 65.60.6 | 97.70.3 | |
@@
| L1-sparse (Jia et al., 2023) | 63.70.1 | 75.20.6 | 56.30.1 | 0.20.1 | 1.30.3 | 69.61.2 | 98.60.2 | |
| SalUn (Fan et al., 2024b) | 63.70.8 | 74.41.5 | 55.50.4 | 0.40.0 | 2.30.3 | 69.01.6 | 98.40.2 | |
| SCRUB (Kurmanji et al., 2024) | 66.70.4 | 80.50.1 | 57.40.5 | 0.40.1 | 1.70.3 | 71.30.4 | 97.40.3 | |
@@
| L1-sparse (Jia et al., 2023) | 65.50.8 | 61.80.3 | 55.50.5 | 0.50.1 | 1.60.8 | 57.00.9 | 98.70.2 | |
| SalUn (Fan et al., 2024b) | 64.50.5 | 60.03.8 | 55.21.0 | 0.50.1 | 1.10.4 | 57.03.8 | 98.00.7 | |
| SCRUB (Kurmanji et al., 2024) | 66.70.6 | 62.41.0 | 55.40.3 | 0.30.1 | 1.70.9 | 55.20.6 | 98.40.2 | |
@@
| L1-sparse (Jia et al., 2023) | 64.30.7 | 66.70.4 | 55.80.4 | 0.40.1 | 2.10.8 | 60.81.0 | 98.30.3 | |
| SalUn (Fan et al., 2024b) | 62.51.6 | 64.62.0 | 54.60.8 | 0.30.1 | 2.20.4 | 60.12.0 | 98.00.3 | |
| SCRUB (Kurmanji et al., 2024) | 65.30.3 | 71.40.6 | 57.00.3 | 0.20.0 | 1.90.9 | 63.70.7 | 97.10.1 | |
@@
B.3 Multi-group unlearning